# Line-of-sight shear in SLACS strong lenses

### v1

@@ 3.3 Shear models
We use three different models for the shear of the strong lensing images. The first is the minimal LOS shear model, described by Equation 7. In this case, the shear model has five parameters: the LOS shear parameters $\gamma_{1}^{\rm LOS}$ and $\gamma_{2}^{\rm LOS}$, the foreground shear parameters $\gamma_{1}^{\rm od}$ and $\gamma_{2}^{\rm od}$ and the LOS rotation parameter $\omega_{\rm LOS}$.
In Hogg et al. (2023), it was shown that when fitting lens models to mock data, the original image could be recovered even when the foreground shear was neglected in $\psi_{\rm eff}$. In the second shear model we consider, we omit the explicit appearance of $\gamma_{\rm od}$ in the effective main-lens potential. The foreground shear term physically corresponds to the effect of post-Born corrections in the main deflection – in the presence of foreground perturbations, light rays do not hit the main lens plane at the same position as if those perturbations were absent. Thus, the omission of the foreground shear must be understood as the assumption that such post-Born corrections are degenerate enough with the parameters of the main lens to be absorbed in them; we do not assume that foreground shear, nor its post-Born effect on the main deflector, are negligible. Practically speaking, this means that we fix $\gamma_{1}^{\rm od}=\gamma_{2}^{\rm od}=0$ in our analysis using this model, which we refer to as the ‘no foreground shear model’ throughout the rest of this paper.
The final shear model we consider is a further restriction of the above, created by fixing the LOS rotation to zero, $\omega_{\rm LOS}=0$. This model is mathematically equivalent to the ‘external’ or ‘residual’ shear model widely used in the literature (Tan et al. (2024); see Shajib et al. (2024) for a discussion of this terminology), and has two parameters, $\gamma_{1}^{\rm LOS}$ and $\gamma_{2}^{\rm LOS}$. We fit the strong lenses with this model to allow us to make a direct comparison with shear values previously reported in the literature for these lenses. We refer to this model as ‘external shear equivalent’.
@@ 3.4 The automated modelling procedure
Before beginning the modelling, we apply a custom annular mask to each of our 50 selected SLACS lenses, with an inner radius of $0.4^{\prime\prime}$ and an outer radius selected manually for each lens (typically around $3.0^{\prime\prime}$) masking out the central and outer regions of the images. We also mask the light from any prominent foreground object in the field of view. This masking is an alternative to the lens light removal, for example via a multi-Gaussian expansion, that is performed in other analyses, e.g. He et al. (2024). However, when the mask is in place, we still model the remaining lens light, allowing us to infer the uncertainties in the model parameters. This cannot be done when the lens light is removed before any modelling step.
@@
We then fit lens models to the images using the lenstronomy package (Birrer & Amara, 2018; Birrer et al., 2021), via the dolphin software (Shajib et al., 2021; Tan et al., 2024), which provides an automated modelling pipeline to uniformly model a large sample of lenses. The modelling has two stages: first, a particle swarm optimisation (PSO) is used to locate the maximum likelihood value for each parameter in the model, i.e. the lens mass, lens light and source light parameters.
@@
The dolphin package provides a recipe to increase the computational efficiency of the optimisation step, in which the lens mass, lens light and source light model parameters are individually optimised whilst the rest of the parameters are kept fixed, before all the model parameters are optimised together. Full details of this optimisation recipe can be found in Shajib et al. (2021).
@@
In the second stage of the parameter inference, an MCMC method is used to sample from the posterior distribution formed by the likelihood and the user-defined priors for each parameter. In this work we use the affine-invariant ensemble sampling algorithm provided by the emcee package (Goodman & Weare, 2010; Foreman-Mackey et al., 2013); the ensemble of walkers is initialised in a small Gaussian ball around the maximum likelihood value found by the PSO. Our chosen priors are listed in Table 1.
Component Parameter Prior
$\theta_{\rm E}$ $[0.03^{\prime\prime},3^{\prime\prime}]$
$\gamma^{\rm EPL}$ $[1.3,2.8]$
Lens mass $e_{1}$ $[-0.5,0.5]$
$e_{2}$ $[-0.5,0.5]$
$x$ $[-0.5^{\prime\prime},0.5^{\prime\prime}]$
$y$ $[-0.5^{\prime\prime},0.5^{\prime\prime}]$
$R_{\rm{eff}}$ $[0.1,5.0]$
$e_{1}$ $[-0.5,0.5]$
Lens light $e_{2}$ $[-0.5,0.5]$
$x$ $[-0.5^{\prime\prime},0.5^{\prime\prime}]$
$y$ $[-0.5^{\prime\prime},0.5^{\prime\prime}]$
$\beta$ $[0.02^{\prime\prime},0.2^{\prime\prime}]$
$R_{\rm{eff}}$ $[0.04^{\prime\prime},0.5^{\prime\prime}]$
$n_{\rm{S}}$ $[0.5,8.0]$
Source light $e_{1}$ $[-0.5,0.5]$
$e_{2}$ $[-0.5,0.5]$
$x$ $[-0.2^{\prime\prime},0.2^{\prime\prime}]$
$y$ $[-0.2^{\prime\prime},0.2^{\prime\prime}]$
$\gamma_{1}^{\rm od}$ $[-0.5,0.5]$
$\gamma_{2}^{\rm od}$ $[-0.5,0.5]$
Shear $\gamma_{1}^{\mathrm{LOS}}$ $[-0.5,0.5]$
$\gamma_{2}^{\mathrm{LOS}}$ $[-0.5,0.5]$
$\omega_{\mathrm{LOS}}$ $[-0.5,0.5]$
@@
However, this is only true for objects with a relatively insignificant foreground; if a non-negligible foreground is present, the lens light itself will undergo potentially significant lensing by that foreground. In this case, a prior rigidly joining lens light and lens mass ellipticity is no longer physically well-motivated, and the degeneracies present between foreground shear, $\gamma_{\rm od}$, and the lens mass ellipticity will lead to biased or simply incorrect parameter inference across the model; see Johnson et al. (2024) for further discussion of foreground biases in strong lensing. We will show an example of this below; in general, we recommend that, given the entanglement between foreground effects and the lens model, this type of prior should not be used when attempting to infer LOS shear in the minimal model (as this model contains $\gamma_{\rm od}$).
The benefits of this kind of automated modelling procedure where the same model is fit to every lens are that the required investigator time is minimised, as it obviates the need for manual modelling of individual lenses to fine-tune the setup (which would anyway be infeasible even for a modest-size lens catalogue, e.g. over 500 lenses), and that results from individual fits are more readily comparable. On the other hand, whilst the strong lensing cross-section is dominated by elliptical galaxies (Turner et al., 1984), there is bound to be a potentially significant number of lenses that cannot be satisfactorily modelled by a ‘one size fits all’ approach. As we will see, this accounts for a significant fraction of the lenses we attempted to model in this work.
@@ 4 Results and discussion
The above-described automated modelling procedure allowed us to successfully model 18 of the initial 50 lenses with the minimal model. We rejected eight fits due to the LOS shear parameters $\gamma_{1}^{\rm LOS}$ and $\gamma_{2}^{\rm LOS}$ being unconstrained, a further thirteen fits due to the presence of strong residuals between the model and the data, and a further three fits due to the best-fit value of the slope of the elliptical power law potential being atypically shallow, $\gamma^{\rm EPL}\leq 1.4$. Note that the remaining eight lenses of the initial 50 were successfully modelled with the no foreground shear or external shear models, but not the minimal model, which we will discuss in more detail below.
Small values of $\gamma^{\rm EPL}$ result in central images in the lens model which would never be observable in reality due to the light emitted by the deflector, but since we mask out the centre of each image during our modelling, the absence of corresponding central images in the data are not penalised (Shajib et al., 2021). We therefore choose to remove the three cases where $\gamma^{\rm EPL}$ has converged to a very small value a posteriori.
4.1 LOS shear in the minimal model
@@
Figure 1: The first six lenses fit successfully with the minimal model. From left to right, the panels show the single-band image data for each lens, our reconstruction of the image along with the reduced $\chi^{2}$ of the model, the residual difference between the image and the reconstruction, the reconstructed source and the joint posterior distribution of the components of the LOS shear, $\gamma_{1}$ and $\gamma_{2}$. The dark and light shaded areas are respectively the $1\sigma$ and $2\sigma$ confidence intervals.
Figure 2: The next six lenses fit successfully with the minimal model. From left to right, the panels show the single-band image data for each lens, our reconstruction of the image along with the reduced $\chi^{2}$ of the model, the residual difference between the image and the reconstruction, the reconstructed source and the joint posterior distribution of the components of the LOS shear, $\gamma_{1}$ and $\gamma_{2}$. The dark and light shaded areas are respectively the $1\sigma$ and $2\sigma$ confidence intervals.
Figure 3: The last six lenses fit successfully with the minimal model. From left to right, the panels show the single-band image data for each lens, our reconstruction of the image along with the reduced $\chi^{2}$ of the model, the residual difference between the image and the reconstruction, the reconstructed source and the joint posterior distribution of the components of the LOS shear, $\gamma_{1}$ and $\gamma_{2}$. The dark and light shaded areas are respectively the $1\sigma$ and $2\sigma$ confidence intervals.
We firstly present the 18 lenses in which the LOS shear was successfully measured using the minimal LOS shear model. The observed image, reconstructed image, normalised residuals between data and model, reconstructed source and joint marginalised posterior distribution of the LOS shear parameters $\gamma_{1}^{\rm LOS}$ and $\gamma_{2}^{\rm LOS}$ are shown in the five respective columns of Figure 1, Figure 2 and Figure 3 for each of the 18 lenses. Furthermore, the reduced $\chi^{2}$ of each model is shown in the upper left part of each panel in the second column, and the measurement of the shear magnitude $\gamma_{\rm LOS}$ and the associated $1\sigma$ uncertainty is shown in each panel of the last column. We use the ChainConsumer package (Hinton, 2016) to analyse our MCMC chains; we present the posterior distributions after the removal of the first 20% of the chain as burn-in.
We note that the reduced $\chi^{2}$ for the majority of our models, i.e. the $\chi^{2}$ normalised by the number of degrees of freedom in the model, is less than one. This usually indicates overfitting; in other words, the number of parameters in the model is so large that the noise is being fit. However, we must be cautious not to over-interpret the reduced $\chi^{2}$ values, as we do not compute the uncertainty on these numbers. For a rough estimation, we may approximate the $\chi^{2}$ distribution as a Gaussian, allowing us to estimate its standard deviation as $\sigma_{\chi^{2}}\approx\sqrt{2/N}$, where $N$ is the number of samples (Andrae et al., 2010). For our 18 successfully fit lenses, $\sigma_{\chi^{2}}\approx\sqrt{2/18}=0.33$, meaning that all but two of our reduced $\chi^{2}$ values are consistent with one at $1\sigma$.
In all of the lenses shown here, the magnitude of the LOS shear is found to be distinct from zero by at least $1\sigma$. We can compare the distribution of the measured LOS shear with ‘external’ or ‘residual’ shear measured in other works. We select the shear measurements in Shajib et al. (2021), Etherington et al. (2022) and Tan et al. (2024), as these three works study similar lens samples to those in this work, using the same or similar modelling pipelines. The samples do not directly overlap; we model eight lenses successfully that were not successfully modelled by Shajib et al. (2021), of which three were also not successfully modelled by Tan et al. (2024): SDSSJ0959+4416, SDSSJ1016+3859 and SDSSJ1020+1122. We successfully modelled five lenses that were not successfully modelled by Etherington et al. (2022); these were all successfully modelled by either Shajib et al. (2021) or Tan et al. (2024).
Figure 4: Histograms of shear magnitude measurements. The three panels in the top row (solid colour histograms) show the external shear as measured in the cited works. The open black histograms show the LOS shear as measured in this work. The bottom row (hatched coloured histograms) show the measurements of external shear from the same lenses as the minimal model was successful at fitting in this work.
In Figure 4 we show histograms of the measurements of the shear magnitude in this work ($\gamma_{\rm LOS}$) and from the aforementioned other works ($\gamma_{\rm ext}$). The top row shows the distributions of all the shear measurements from those works, whilst the bottom row shows the distributions of the shear measurements from just those lenses which we modelled successfully with the minimal LOS shear model in this work.
We can see from this figure that when comparing to the full distributions of shear measurements (solid colour histograms), the distribution of the measured LOS shear magnitude (black) is somewhat broader than the measurements of the external shear magnitude from the three works we compare with. The LOS shear distribution has a longer tail reaching to larger shear magnitudes, except for one outlying point in the Etherington et al. (2022) measurements. Furthermore, the peak of the LOS shear distribution is at a larger magnitude than the external shear from the works we compare with. The same holds true for the comparison with the shears from only those lenses which match those modelled by the minimal model in this work (coloured hatched histograms).
To quantify the differences in the distributions, we can use the $k$-sample Anderson–Darling (AD) test (Anderson & Darling, 1952), which tests whether a number of different samples are drawn from the same underlying population. The test is similar to the Kolmogorov–Smirnov test (Sprent, 1998), but is more sensitive to differences in the tails of the sample distributions (Razali & Wah, 2011). This is appropriate for our distributions of measured shear. We run the AD test three times, comparing the distribution of the LOS shear magnitudes with the distribution of the external shear magnitudes in Shajib et al. (2021), Etherington et al. (2022) and Tan et al. (2024). We use only the shears as measured from the same lenses modelled in this work.
At 5% significance level, i.e. approximately $2\sigma$, the null hypothesis that the two samples are drawn from the same underlying distribution is rejected for the shears measured in Shajib et al. (2021), but cannot be rejected in the other two cases. From this, we conclude that the information contained in measurements of the shear parameter in the minimal model does not differ from that contained in the external shear to a statistically significant level, in the lenses studied in this work.
4.2 Neglecting foreground shear and LOS rotation
Given this finding, the role of the foreground shear $\gamma_{\rm od}$ and the LOS rotation, $\omega_{\rm LOS}$, may be questioned. If the distribution of LOS shear is statistically indistinguishable from distributions of external shear found in other studies, what value do the additional three parameters bring? We can explore this question by re-fitting the same strong lenses but with the foreground shear fixed to zero, and again with both the foreground shear and LOS rotation fixed to zero. The latter model is mathematically equivalent to the external shear model, as it contains only two parameters to describe the shear effect, $\gamma_{1}^{\rm LOS}$ and $\gamma_{2}^{\rm LOS}$.
Our first finding is that seven more lenses on top of the 18 fit with the minimal model were able to be modelled successfully in the no foreground shear model, and four more lenses with the external shear equivalent model, of which three lenses are common to both sets. Six of these fits were rejected in the case of the minimal model due to the LOS shear components being unconstrained; the remaining two had very small best-fit values of the EPL slope. This finding implies that, as is commonly known, the reduction in parameter space may bring a marginal improvement in the sampling of the MCMC; fewer parameters means that the joint posterior distribution is simpler and potentially easier to sample from, leading to better convergence.
Figure 5: Measured LOS shear in the minimal model (solid blue), no foreground shear model (unfilled light blue) and external shear equivalent model (unfilled dark blue).
In Figure 5 we show the distributions of the measured LOS shear in the minimal model, no foreground shear and external shear equivalent models. The distributions are broadly similar in shape, implying that the reduction in parameter space does not significantly change the measured LOS shear distribution, as expected from the comparison between our results and the literature. The two-sample AD test does not reject the null hypothesis that the samples were drawn from the same underlying distribution as the minimal model in either case.
@@ 4.2 Neglecting foreground shear and LOS rotation
Figure 7: Left: LOS shear magnitudes in the minimal model plotted against the LOS shear magnitudes in the no foreground shear model (blue crosses). Right: LOS shear magnitudes in the minimal model plotted against the LOS shear magnitudes in the external shear equivalent model (blue crosses). The dashed grey lines represent the $y=x$ line whilst the solid blue lines show a linear fit to the measurements, with a $1\sigma$ shaded region. The fits are $y=0.76x+0.02$ and $y=1.23x-0.01$ respectively. The error bars on the points are the $1\sigma$ standard deviations.
Next, we can examine the posterior distributions of the lens ellipticity and shear parameters to understand how the degeneracies between these quantities behave in each model. As an example, in Figure 6, we show the marginalised posterior distributions on the lens mass ellipticity, foreground shear, LOS shear and LOS rotation in the minimal model, no foreground shear model and external shear equivalent model as fit to lens SDSS03300-0020.
From this figure, the extremely strong degeneracy between the lens mass ellipticity and foreground shear can be seen in the minimal model (blue contours). Removal of the foreground shear from the model (purple contours) leads to spuriously tight constraints on the ellipticity of the lens mass. The degeneracy between the foreground shear and the lens mass ellipticity is propagated to the LOS shear, leading to a significantly different LOS shear measurement. In this lens, the degeneracy is particularly apparent between $\gamma_{1}^{\rm LOS}$ and $e_{1}^{\rm mass}$, whilst the central values of the $\gamma_{2}^{\rm LOS}$ posteriors change significantly between the minimal and no foreground shear models.
This finding is in sharp contrast with the results of Hogg et al. (2023), which demonstrated that, in mock data, neglecting the foreground shear in the model had little effect on the inferred LOS shear. However, we are now considering a far more complex model applied to real data. Somewhat surprisingly, when the LOS rotation is removed from the model (green contours), the LOS shear constraint is once again in agreement with the minimal model.
To examine whether this effect is systematic across all the lenses studied in this work, we can look at the LOS shear magnitudes measured in the minimal model compared to those measured in the no foreground shear and external shear equivalent models. In Figure 7 we show these measurements plotted against each other; the minimal model against the no foreground shear model in the left panel and the minimal model against the external shear equivalent model in the right panel.
Computing a linear fit [^2] To do this, we use the ordinary least squares fitting method provided by numpy.polyfit. However, least squares fitting becomes inaccurate when the independent variable has significant uncertainties, as is the case here (Fuller, 1987). We further choose not to weight the fit using the uncertainty on the dependent variable in either case. The displayed uncertainty on each linear fit (shaded blue region) is therefore computed from the covariance matrix returned by numpy.polyfit based solely on the shear values, without their errors. to the scatter shows that, for large shear values ($\gamma_{\rm LOS}>0.1$), the LOS shear magnitude measured in the no foreground shear model tends to be slightly smaller than what would be measured in the minimal model, whilst for small values of the shear, the results from the two models are consistent. Conversely, the LOS shear magnitude measured in the external shear model tends to be larger than what would be expected from the minimal model, again for large shear values.
The slope of the fit being less than one in the comparison between the minimal model and no foreground shear model is driven by the one clear outlying point at $\gamma_{\rm LOS}=0.25$ in the minimal model and $\gamma_{\rm LOS}=0.07$ in the no foreground shear model. Removing this point results instead in a slope greater than one, similar to that seen when comparing the minimal model with the external shear model.
Furthermore, this figure reveals that the uncertainty on the measured LOS shear magnitude tends to increase as the magnitude itself increases. Due to the small number of points and their relatively large uncertainties, we note that this conclusion may not be borne out if a larger number of more accurate measurements were to be achieved.
We may conclude here that, when considering a single strong lens, a measurement of LOS shear has the potential be significantly affected by neglecting foreground shear and rotation, and the subsequent inheritance of a degeneracy between the LOS shear parameters and the lens mass ellipticity parameters. The difference between the models is negligible for small inferred shear values but grows with the magnitude of the shear. It is therefore apposite to investigate possible causes of these large shear values.
4.3 Adding boxyness and diskyness
Etherington et al. (2024) showed that the large values of shear measured from strong lensing are inconsistent with expectations from weak lensing measurements, both in mock data and in observations, a study which was conducted using the external shear model. We are able to make an independent confirmation of this finding for LOS shear using $N$-body simulations.
We use the RayGalGroupSims suite of relativistic $N$-body simulations, whose publicly available data contains weak lensing maps that include post-Born effects, magnification bias and redshift-space distortions (Rasera et al., 2022). The maps provide convergence and shear measurements at a given set of redshifts, between which we can interpolate to compute $\gamma_{\rm os}$, $\gamma_{\rm od}$ and $\gamma_{\rm ds}$ shear values for any given deflector and source redshift. From these, we compute the expected $\gamma_{\rm LOS}$ for any line of sight in the simulation, and thus produce a distribution of expected $\gamma_{\rm LOS}$ shear magnitudes.
In Figure 8, we show the distribution of the two LOS shear components $\gamma_{1}^{\rm LOS}$ and $\gamma_{2}^{\rm LOS}$, as well as the magnitude of the LOS shear $\gamma_{\rm LOS}$ as estimated from the RayGalGroupSims (unfilled black histograms) along with the measurements from the SLACS lenses (solid blue histograms). Looking first at the distributions of the two individual shear components, we do not see a stark difference between the simulated and observed shears; in particular, the peaks of the distributions are in good agreement. There remains a clear difference in the tails of the simulated and observed distributions of both shear components.
Examining instead the shear magnitude, the significant difference between the shear estimated from the simulations and the shear observed in the strong lenses is clearly apparent, in agreement with the findings of Etherington et al. (2024). The two-sample AD test rejects the null hypothesis that the two samples (simulated and observed) are drawn from the same underlying distribution at $95\%$ confidence.
Figure 8: Distribution of the LOS shear measurements from the SLACS lenses using the minimal model (solid blue) and as estimated from the RayGalGroupSims simulations (unfilled black). The left panel shows the $\gamma_{1}$ component of the LOS shear, the middle panel shows the $\gamma_{2}$ component and the right panel shows the LOS shear magnitude.
It therefore seems possible that there is some contribution to the LOS shear from the local environment of the lens galaxies, for example produced by sub-haloes within the main dark matter halo of the deflector. If present, these sub-haloes may also produce non-quadrupolar distortions to the lens image. Furthermore, there is evidence from observations (Bender et al., 1988; Stacey et al., 2024) and simulations (Naab et al., 1999) that isophotes of elliptical galaxies show octupolar distortions (so-called ‘boxyness’ and ‘diskyness’), which, it could be argued, should also be reflected in their underlying mass distribution.
One conclusion that may be drawn from this is that ‘external’ shears may be biased high due to unaccounted-for boxyness or diskyness in the deflector mass profile. We can test whether this conclusion holds for the LOS shear by adding boxyness and diskyness to our lens model and re-fitting the same lenses as already modelled with the minimal model.

### v2

@@ 3.3 Shear models
We use two different models for the shear of the strong lensing images. The first is the minimal LOS shear model, described by Equation 8. In this case, the shear model has five parameters: the LOS shear parameters $\gamma_{1}^{\rm LOS}$ and $\gamma_{2}^{\rm LOS}$, the foreground shear parameters $\gamma_{1}^{\rm od}$ and $\gamma_{2}^{\rm od}$ and the LOS rotation parameter $\omega_{\rm LOS}$.
In Hogg:2022ycw, it was shown that when fitting lens models to mock data, the original image could be recovered even when the foreground shear was neglected in $\psi_{\rm eff}$. In the second shear model we consider, we omit the explicit appearance of $\gamma_{\rm od}$ in the effective main-lens potential. The foreground shear term physically corresponds to the effect of post-Born corrections in the main deflection – in the presence of foreground perturbations, light rays do not trace to the same position in the observer’s plane as they would if those perturbations were absent.
Thus, the omission of the foreground shear must be understood as the assumption that such post-Born corrections are degenerate enough with the parameters of the main lens to be absorbed in them; we do not assume that foreground shear, nor its post-Born effect on the main deflector, are negligible. Practically speaking, this means that we fix $\gamma_{1}^{\rm od}=\gamma_{2}^{\rm od}=0$ in our analysis using this model, which we refer to as the ‘no foreground shear model’ throughout the rest of this paper.
These models are both mathematically distinct from the notion of ‘external’ or ‘residual’ shear which is widely used in the literature (Tan2024; see Shajib2024 for a discussion of this terminology). The LOS formalism can be reduced to the external shear model by fixing both the foreground shear and LOS rotation to zero, $\gamma_{\rm od}=\omega_{\rm LOS}=0$, which leaves the shear to be described by just two complex parameters, $\gamma_{1}$ and $\gamma_{2}$.
@@ 3.4 The automated modelling procedure
Before beginning the modelling, we apply a custom annular mask to each of our 23 selected SLACS lenses, with the inner and outer radii selected manually for each lens (typically around $0.4^{\prime\prime}$ and $3.0^{\prime\prime}$ respectively) masking out the central and outer regions of the images. We also mask the light from any prominent foreground object in the field of view. This masking is an alternative to the lens light removal, for example via a multi-Gaussian expansion, that is performed in other analyses, e.g. He:2024udi. However, when the mask is in place, we still model the remaining lens light, allowing us to infer the uncertainties in the model parameters. This could not be done if the lens light is removed before any modelling step.
@@
We then fit lens models to the images using the lenstronomy package (Birrer:2018xgm; Birrer2021), via the dolphin software (Shajib2021; Tan2024; Shajib:2025bho), which provides an automated modelling pipeline to uniformly model a large sample of lenses. The modelling has two stages: first, a particle swarm optimisation (PSO) (Eberhart1995) is used to locate the maximum likelihood value for each parameter in the model, i.e. the lens mass, lens light and source light parameters.
@@
The dolphin package provides a recipe to increase the computational efficiency of the optimisation step, in which the lens mass, lens light and source light model parameters are individually optimised whilst the rest of the parameters are kept fixed, before all the model parameters are optimised together. Full details of this optimisation recipe can be found in Shajib2021.
@@
In the second stage of the parameter inference, an MCMC method is used to sample from the posterior distribution formed by the likelihood and the user-defined priors for each parameter. In this work we use the affine-invariant ensemble sampling algorithm provided by the emcee package (GoodmanWeare; ForemanMackey2013); the ensemble of walkers is initialised in a small Gaussian ball around the maximum likelihood value found by the PSO. Our chosen priors are listed in Table 2.
The positions of the walkers in the ensemble are updated using a random choice of either the differential evolution (DE) (Nelson2014) and DE ‘snooker’ moves (terBraak2008), with a respective probability of 80% and 20%, which is more efficient [^2] See the emcee documentation: https://emcee.readthedocs.io/en/stable/tutorials/moves. for sampling from mildly multimodal distributions than the standard ‘stretch’ move of GoodmanWeare.
Multimodality may be encountered in our inference when foreground shear and the ellipticity of the EPL profile conspire to produce a given image with interchangeable orientation angles; for example, the combination of $|\gamma_{\rm od}|$ at an angle of $0\degree$ East of North combined with an EPL ellipticity of $|e_{\rm EPL}|\equiv|\gamma_{\rm od}|$ at $45\degree$ East of North is indistinguishable in effect from the same quantities with their angles exchanged. This can result in bimodal marginalised posterior distributions for the components of these complex variables. The effect is simply a manifestation of the constructed degeneracy between $\gamma_{\rm od}$ and lens mass ellipticity in the minimal LOS model.
@@
However, this is only true for objects with a relatively insignificant foreground; if a non-negligible foreground is present, the apparent orientation of the lens light will generally differ from both the apparent and intrinsic orientation of the mass model ellipticity. In this case, a prior rigidly joining lens light and lens mass ellipticity is no longer physically well-motivated, and the degeneracies present between foreground shear, $\gamma_{\rm od}$, and the lens mass ellipticity will lead to biased or simply incorrect parameter inference across the model; see Johnson:2024hvl for further discussion of foreground biases in strong lensing. Given the entanglement between foreground effects and the lens model, this type of prior should not be used when attempting to infer LOS shear.
@@ 4 Results and discussion
The above-described automated modelling procedure allowed us to successfully model all 23 lenses in the Shajib2021 sample. Our results are presented after post-processing the chains, including the removal of burn-in and smoothing of the marginalised posterior distribution histograms.
4.1 Shear is measurable in the minimal LOS model
In Figure 1, Figure 2, Figure 3, and Figure 4, we show the results of our lens modelling with the minimal LOS shear model, and the other model components described above. The first column of these figures shows the image data for each strong lens; the second column shows the model reconstruction of the image and associated $\chi^{2}$ value; the third column shows the residual difference between the data and model; the fourth column shows the reconstruction of the source; and the fifth and final column shows the 1D marginalised posterior distribution of the LOS shear magnitude, $|\gamma_{\rm LOS}|$, for that system. The posterior samples for this quantity are computed directly from the posterior samples of the individual shear components, $\gamma_{1}$ and $\gamma_{2}$.
@@ 4.1 Shear is measurable in the minimal LOS model
Figure 1: The first six lenses fit with the minimal model. From left to right, the panels show the single-band image data for each lens, our reconstruction of the image along with the reduced $\chi^{2}$ of the model, the residual difference between the image and the reconstruction, the reconstructed source and the one dimensional marginalised posterior distribution of the LOS shear magnitude, $|\gamma_{\rm LOS}|$. The shaded area is the $1\sigma$ confidence interval.
Figure 2: The next six lenses fit with the minimal model. The panels show the same information as in Figure 1.
Figure 3: The next six lenses fit with the minimal model. The panels show the same information as in Figure 1.
Figure 4: The final five lenses fit successfully with the minimal model. The panels show the same information as in Figure 1.
In Figure 5, we show the histogram of the magnitude of the LOS shears measured in our sample of 23 lenses in solid blue, compared with the histogram of the ‘external’ shear found in the same lenses by Shajib2021. From this figure, we can see that our shear measurements are very consistent with the previous study by Shajib2021, with the exception of the single outlier SDSSJ1531-0105 at $|\gamma_{\rm LOS}|=0.22\pm 0.03$.
Figure 5: This figure shows the histograms of the $|\gamma_{\rm LOS}|$ values measured in this work (solid blue) and the external shear values measured in Shajib2021 (unfilled black).
4.2 The measured shear is larger than expected
Across the set of 23 lenses, we measure the LOS shear with a mean magnitude of $0.056\pm 0.016$. Etherington:2023yyh showed that the large values of external shear measured from strong lensing are inconsistent with expectations from weak lensing measurements, both in mock data and in observations of $\gamma_{\mathrm{os}}$ in the same regions of the sky as the strong lenses, a study which was conducted using a sample of the SLACS lenses which overlaps with ours (13 lenses in common). We are able to make an independent confirmation of this finding for LOS shear using $N$-body simulations.
We use the RayGalGroup suite of relativistic $N$-body simulations, whose publicly available data contains weak lensing maps that include post-Born effects, magnification bias and redshift-space distortions (2019MNRAS.483.2671B; Rasera:2021mvk). The maps provide convergence and shear measurements at a given set of redshifts, between which we can interpolate to compute $\gamma_{\rm os}$, $\gamma_{\rm od}$ and $\gamma_{\rm ds}$ shear values for any given deflector and source redshift, following the method presented in Johnson2025. From these, we compute the expected $|\gamma_{\rm LOS}|$ for any line of sight in the simulation, and thus produce a distribution of expected $|\gamma_{\rm LOS}|$ shear magnitudes for either an ensemble of lines of sight, or for specific ($z_{\rm d}$, $z_{\rm s}$) pairs.
Figure 6: This figure shows the measured value of $|\gamma_{\rm LOS}|$ and the $1\sigma$ error bars for each lens in our sample. The black dotted line shows the mean expected value of $|\gamma_{\rm LOS}|$ for that system’s lens and source redshift, computed from the RayGalGroup simulations, as described in the main text. The grey shaded regions represent the $1$, 33 and $5\sigma$ uncertainties around this mean. The measured values are coloured by their tension in $\sigma$ from the mean, defined in Equation 14.
In Figure 6, we show our measured LOS shear value in each system against the mean expected $|\gamma_{\rm LOS}|$ for the deflector and source redshift of that observed system, obtained from the simulations. The black dotted line shows the mean, whilst the grey bands represent the $1$, 33 and $5\sigma$ uncertainties. The measured points are coloured according to their tension with the simulation. The tension between two quantities $a$ and $b$ is defined as follows,
@@ 4.2 The measured shear is larger than expected
where $x$ are the best-fit values and $\sigma$ the standard deviations of the two quantities being compared.
From Figure 6, we can see that the majority of measured $|\gamma_{\rm LOS}|$ values are larger than the associated expected mean, with the largest tension between measurement and simulation sitting at $7.9\sigma$ for lens SDSSJ1531+0105 (all numerical tension values are listed in Table 1). A further six lenses have tensions of $3\sigma$ or more with the expectation from the simulation. In the next subsection, we will explore one possible explanation for these unexpectedly large shear values.
4.3 Shear does not compensate for unmodelled octupoles
Previous works have asserted that the large values of external shear measured in strong lensing images are the result of unmodelled complexity in the dominant lens, proposing that a fourth-order multipole, or octupole, in the lensing potential may be responsible (Etherington:2023yyh). This so-called ‘boxyness’ or ‘diskyness’ has also been observed in the isophotes of elliptical galaxies (Bender1988; Stacey2024) and produced in simulations (Naab:1999ge). In this subsection, we investigate the effect including octupoles in our lens model has on the inferred LOS shear.